### v4

@@ 3.1 Synthetic data generation and Gibbs implementation
[…]
    $\begin{matrix}X_{1i}\sim\mathrm{N}(200,100^{2}),{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}...........}\\ X_{2i}\sim\mathrm{N}(100,100^{2}),{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}...........}\\ X_{3i}\sim\mathrm{N}(20,50^{2}),{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}..............}\\ X_{4i}\sim\mathrm{N}(50,50^{2}),{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}..............}\\ X_{5i}\sim\mathrm{N}(2,5^{2}){\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}..............}\text{and}\\ X_{6i}\sim\mathrm{N}(25,50^{2})\text{;}\qquad\qquad{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}.}\end{matrix}$
@@
    Define the functional coefficients
    $\begin{matrix}\beta_{0}(t)=\exp(t^{2})\text{,}\qquad{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}....}\\ \beta_{3}(t)=\cos(2t)\qquad\text{and}\\ \beta_{5}(t)=t^{3}\text{,}\qquad\qquad\quad{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}...}\end{matrix}$
    in which $t\in[0,2]$;
@@
    All other partial functional coefficients do not need to be defined since it is considered that only two covariates $X_{3i}$ and $X_{5i}$, should be used for data generation;
@@
The equality presented in 13 is a summary representation, since its full version is given by
@@
The covariates chosen to compose the discrete representation (13) were purposely positioned in the design matrix $\boldsymbol{X}$ so that both were not neighbors. Thus, we can verify that the model truly has the ability to select covariates and not just blocks of covariates.
Therefore, any method that performs a good fit to this dataset should be able to identify that only $X_{3i}$ and $X_{5i}$ are useful covariates, in addition to being able to calculate estimates for the functional coefficients which are close to the original functional parameters. When building the synthetic data, we establish that $\beta_{3}(t)$ and $\beta_{5}(t)$ are the so-called “functional coefficients” instead of “partial functional coefficients” since it is implied that the latent variable is equal to one in the generation of synthetic data in both cases.
@@
Bayesian inference was implemented through the Gibbs sampler described in Section 2.2 using two chains to enable the convergence diagnosis, which start at different points, and with the calculation of 10 000 iterations. Considering a burn-in period of 50% of the size of each chain and spacing of 50 points between each sampled value, each chain contains 100 sampled points. Thus, at the end of the process, a posterior sample of size 200 will be obtained for each parameter and latent variable.
The convergence diagnosis used in our analyses takes into account the convergence test of the chains proposed by [12]. In our numerical experiments carried out herein, the chains were initialized as follows:
@@
The latent random variables $Z_{l}$’s were initialized in the first chain via random draw, and the complementary vector of the drawn variables for the first chain was used for the second chain.
With regard to the hyperparameters, it was defined: $\delta_{1}=\delta_{2}=0$. Thus, when using a degenerated prior to $\sigma^{2}$ with $\delta_{1}=0$ and $\delta_{2}=0$, there is an equivalence with the use of a non-informative improper prior $\frac{1}{\sigma^{2}}$. Then for the hyperparameter of the priors of the $\tau_{kl}$’s, $\lambda=\sqrt{2}$ was defined such that the prior mean is equal to 1. This value of $\lambda$ is also the smallest possible value (in this data set), since uniqueness problems in the stationary distribution begin to appear for smaller values of this parameter as a result of the combination of two factors: low regularization and multicollinearity between the covariates. Therefore, there was a configuration with the lowest possible regularization.
In situations where $\boldsymbol{\mu}$ is considered as a parameter, $\psi=0.6$ was defined and the first chain was initialized with $\boldsymbol{\mu}=\boldsymbol{\frac{1}{5}}$, while the second was with $\boldsymbol{\mu}=\boldsymbol{\frac{4}{5}}$. On the other hand, several models were generated for cases in which such a component was considered as a hyperparameter, taking into account a grid of nine equally spaced elements ranging from $0.1$ to $0.9$.
To summarize the resulting information, we calculated, for each parameter, the MAP estimate (Maximum a Posteriori). Thus, regarding the latent variables $Z_{l}$’s, we obtained the most frequent value, meaning the posterior mode.
@@ 3.3 Results
This section exhibits the results of two different types of numerical experiments. Firstly, the proposed method is tested with synthetic data without replications. Secondly, simulations show the performance of the procedure under a scheme with replications. The diagnostic analysis based on the method proposed by [12] attested the convergence of the chains of the partial functional coefficients in all experiments and in all model configurations evaluated after the burn-in period. It is also important to notice that B-splines were used for the basis expansion of all partial functional coefficients.
3.3.1 Studies with Synthetic Data
[…]
$\mu$ (Hyperparameter)
$\sigma$ $m$ $K$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 $\mu$ (Parameter)
5 0.99945 0.99946 0.99946 0.99947 0.99945 0.99945 0.99946 0.99945 0.99945 0.99947
10 0.99993 0.99989 0.99987 0.99989 0.99990 0.99988 0.99991 0.99990 0.99992 0.99985
10 15 0.99968 0.99976 0.99974 0.99951 0.99970 0.99971 0.99970 0.99965 0.99963 0.99957
5 0.99920 0.99920 0.99920 0.99920 0.99920 0.99919 0.99919 0.99919 0.99919 0.99919
10 0.99995 0.99994 0.99994 0.99995 0.99996 0.99995 0.99995 0.99995 0.99995 0.99995
0.2 15 15 0.99990 0.99990 0.99983 0.99986 0.99988 0.99991 0.99983 0.99990 0.99992 0.99990
5 0.81410 0.81418 0.81381 0.81367 0.81343 0.81447 0.81391 0.81222 0.81485 0.81333
10 0.80982 0.81346 0.81047 0.81186 0.81342 0.81209 0.80735 0.81158 0.81121 0.81316
10 15 0.80450 0.81138 0.81028 0.80865 0.80474 0.81268 0.80949 0.80975 0.80985 0.81112
5 0.75635 0.75703 0.75558 0.76183 0.76300 0.76284 0.76297 0.76271 0.76340 0.75593
10 0.75773 0.75707 0.75726 0.75795 0.75770 0.75702 0.75774 0.75737 0.76165 0.75748
20 15 15 0.75626 0.75634 0.75537 0.75601 0.75300 0.75771 0.75456 0.75414 0.75423 0.75782
Table 1: Performance metrics (Eq. 15) of the proposed Bayesian approach with $\mu$ as a hyperparameter and as a parameter, according to the different configurations of $K$ B-spline basis functions, two dataset lengths ($m\in\{10,15\}$) and two noise variances, $\sigma^{2}=(0.2)^{2}$ and $\sigma^{2}=(20)^{2}$.
Table 1 presents the performance metrics calculated following the Expression (15) for the various proposed model configurations. The highlighted cells in the table indicate the best model configuration among those whose $\mu$ is taken as a hyperparameter. According to the metric (15), the model explains almost all of the variability of the data generated with $\sigma=0.2$. At the same time, it can explain more than 75% of the total variability when $\sigma=20$. A relevant aspect is that the model configurations considering $\mu$ as a parameter obtained very similar performance to the cases with $\mu$ as a hyperparameter.
When assessing the results in Table 1, there is evidence that the proposed model has a low sensitivity in relation to the possible values that the hyperparameter $\mu$ can assume, regardless of the degree of data dispersion ($\sigma$), the number of curves ($m$) and the number of bases ($K$) used in the expansion of the partial functional coefficients. This low sensitivity to the hyperparameter value is a desired property, but it should be checked with more caution in a study with replications, as it is presented in Section 3.3.2.
Supplementary Table 1 displays the Mean Squared Error (MSE) as in (16) of the respective model settings. The best model for small variance ($\sigma=0.2$) gives an MSE of the order $10^{-2}$, while the MSE for $\sigma=20$ is of the order $10^{-1}$. As expected, the ability of the model to estimate the mean functional response in the study with a large $\sigma$ is lower than in the study with a small $\sigma$. Finally, by comparing the MSEs it is possible to identify that the model performed predominantly better with data with $m=15$ than with data with ten curves ($m=10$).
@@ 3.3.1 Studies with Synthetic Data
As an example, Figure 1 presents the estimates $\hat{\beta}_{l}(.)$’s of the partial functional coefficients selected by the model (data with $m=10$ and configurations with $K=10$ and $\mu=0.1$ for the two dispersion levels) and the estimates of the functional intercept $\hat{\beta}_{0}(.)$. Figure 1 also shows the 95% credible bands for the functional coefficients the model selected. We computed these bands as follows. First, the two hundred vectors of parameters sampled by the Gibbs sampler were taken to construct the bands. Then the curves of the functional coefficients were generated through each of these vectors, totalling two hundred curves per functional coefficient. With this, by fixing $t$, it is possible to calculate the necessary quantiles for the construction of the credible bands since there are two hundred possible points for each functional coefficient and each $t$. The quantiles used to define the lower and upper bands were $2.5\%$ and $97.5\%$.
The first relevant fact to be highlighted in Figure 1 is the ability of the model to exactly select the covariates that were used in the generation of synthetic data when using the configuration with $K=10$ and $\mu=0.1$. Observing the posterior samples of each $Z_{l}$ (not shown here), it is also verified that there was perfect agreement between the sampling points and the real values of the latent variables, meaning that a sample was obtained with all quantities equal to one for the cases in which $Z_{l}=1$ in the synthetic data, as well as a sample with all quantities equal to zero for the cases in which $Z_{l}=0$. This is seen in both data dispersion scenarios ($\sigma=0.2$ and $\sigma=20$).
The model’s ability to select the covariates used in the synthetic data generation remains high with other $K$ and $\mu$ configurations. This property will be further investigated in the study with replications presented in Section 3.3.2. A high fitting capacity is also observed in Figure 1, mainly in the scenario with low dispersion data. The true and estimated curves overlap, making it impossible to distinguish them only by graphic visualization. The estimation naturally loses precision when considering data with a higher noise level, but the model still provides satisfactory estimates.
(a) Data with $\sigma=0.2$.
(b) Data with $\sigma=20$.
Figure 2: Partial estimated functional coefficients that were excluded by the model, according to the data dispersion degree.
Although only the functional coefficients $Z_{3}\beta_{3}(t)$ and $Z_{5}\beta_{5}(t)$ are different from zero, since $\hat{Z}_{3}$ and $\hat{Z}_{5}$ are equal to one, the proposed model still provides the estimates of the other $\beta_{l}(t)$’s, even if the latent variables nullify such estimates. Figure 2 presents the estimates of all other partial functional coefficients associated with the covariates excluded from the model. All the functional estimates in Figure 2 are around zero, regardless of the degree of dispersion of the synthetic data used. This pattern for the partial functional coefficients of the excluded covariates is expected due to the very characterization of the full conditional distributions of the basis expansion coefficients (the $b$’s) as a function of the respective latent variables.
@@
To conclude the study of this section, Figure 3 presents the first three mean (or expected) functional response curves (among the total of $m=10$), together with the respective observed data and the respective estimated curves for each data dispersion level. In addition, credible bands for the respective mean functional responses are also presented. The bands are generated similarly to those obtained for the functional coefficients in Figure 1. More specifically, from the two hundred parameter vectors sampled by the Gibbs sampler, the estimated mean functional response curves corresponding to each of those vectors were generated, totalling two hundred curves per functional response. From there, the $2.5\%$ and $97.5\%$ quantiles are calculated as previously described and used to define the lower and upper bands.
3.3.2 Simulations
In this section, we conducted a study with replications to obtain more consistent conclusions regarding our proposed methodology’s fit performance and model selection capacity. In addition, the sensitivity of the proposed model to the hyperparameter $\mu$ is also investigated. The synthetic data of each replication contains $m=10$ curves generated as described in Section 3.1 with $K=10$ B-spline basis functions considered for the expansion of the functional coefficients.
We conducted our simulations considering two dispersion levels for the data ($\sigma=0.2$ and $\sigma=20$). Thus, ten model configurations were tested under each dispersion scenario, nine of which consider $\mu$ as a hyperparameter and one that considers it as a parameter. In all model configurations, regardless of the dispersion scenario, a total of one hundred replications were used.
(a) Data with $\sigma=0.2$.
(b) Data with $\sigma=20$.
Figure 4: Boxplots of the metric (15) for the various model configurations tested when $\mu$ is a hyperparameter, according to the data dispersion degree.
Figure 4 shows the boxplots for metric (15) for each tested model configuration in our simulations. It is noticed that the proposed model, regardless of the dispersion scenario, has a low sensitivity to the hyperparameter $\mu$. This property guarantees that the model can smooth and select the partial functional coefficients (and consequently the covariates), regardless of the prior choice for $\mu$. Supplementary Figure 1 presents the MSE boxplots for each tested model configuration. It is also possible to conclude the low sensitivity of the proposed model to the choice of the hyperparameter $\mu$. Both metrics show an expected behavior concerning $\sigma$ since the performance of the models is better when the data present low dispersion ($\sigma=0.2$). It is also worth mentioning that both metrics are well-behaved because there is symmetry and a few discrepant values in the boxplots.
We also conducted simulations to evaluate the model’s performance considering $\mu$ as a parameter. Supplementary Figure 2 shows the boxplots of the metric (15) according to the data dispersion level. Similarly, Supplementary Figure 3 presents the boxplots of the MSEs. For both data dispersion levels and regardless of the metric being analyzed, it is clear that the model configuration with $\mu$ as a parameter is as efficient as the other configurations in terms of goodness of fit and estimation quality.
In addition to the goodness of fit, it is also interesting to assess the variable selection ability for each model configuration among the one hundred replications. Table 2 presents the proportion among the 100 replications that each functional coefficient estimate ($\hat{Z}_{l}\hat{\beta}_{l}(.)$, for $l=1,\ldots,6$) was different from zero. For data with a low dispersion level ($\sigma=0.2$), we can observe in Table 2 that all tested model configurations correctly performed the selection in all replications. However, the selection for the data with a higher dispersion level ($\sigma=20$) is not perfect in all replications, although it remains highly accurate.
@@ 3.3.2 Simulations
Table 2 shows that except for configurations with $\mu$ as a parameter and as hyperparameter with $\mu=0.1$ when $\sigma=20$, all other configurations (for both $\sigma=0.2$ and $\sigma=20$) correctly selected the covariates associated with the partial functional coefficients $\hat{\beta}_{3}(.)$ and $\hat{\beta}_{5}(.)$ in all replications, allowing us to conclude that the model has no difficulty in selecting the partial functional coefficients that should remain in the functional regression. Furthermore, we can see that most of the selection errors committed in the scenario with $\sigma=20$ happened when the model fails to exclude a coefficient that should be excluded. The proportions associated with the coefficients that should be excluded from the regression are generally below 7%. However, it is worth noting that slightly higher proportions were observed for the configuration that considers $\mu=0.9$.
@@
Similarly, Supplementary Figure 4 shows the estimated partial functional coefficients ($\hat{\beta}_{3}(.)$ and $\hat{\beta}_{5}(.)$) and functional intercept for the case when the model considers $\mu$ as a parameter. The conclusions are similar to the ones obtained from Figure 5, showing that the version with more layers in the hierarchical model does not significantly differ from the version which considers $\mu$ fixed.
(a) Data with $\sigma=0.2$ ($\mu=0.3$).
(b) Data with $\sigma=20$ ($\mu=0.5$).
Figure 6: Estimated partial functional coefficients ($1^{\text{st}}$, $2^{\text{nd}}$, $4^{\text{th}}$ and $6^{\text{th}}$) from the model that considers $\mu$ as a hyperparameter, according to the data dispersion degree, for the replications whose coefficients were excluded among the total of 100 replications.
To complement the analysis of the proposed methodology through simulations, Figure 6 presents the estimates of all the other partial functional coefficients ($\hat{\beta}_{1}(.)$, $\hat{\beta}_{2}(.)$, $\hat{\beta}_{4}(.)$ and $\hat{\beta}_{6}(.)$) from the model that considers $\mu$ as a hyperparameter that are associated with the excluded variables, according to the data dispersion degree (Supplementary Figure 5 presents the corresponding plots obtained by the model that considers $\mu$ as a parameter). Regardless of the data dispersion degree and the approach to $\mu$, Figure 6 and Supplementary Figure 5 shows that all the functional estimates in question are around zero.
3.4 grLASSO, grSCAD and grMCP versus proposed model

### v5

@@ 3.1 Simulated data generation and Gibbs implementation
    Six covariates are generated ($X_{li}$’s, in which $i\in\{1,\;2,\dots,\;m\}$ and $l\in\{1,\;2,\dots,\;p=6\}$) represented by six vectors of $m$ random variables such that $X_{1i}\sim\mathrm{N}(200,100^{2})$; $X_{2i}\sim\mathrm{N}(100,100^{2})$; $X_{3i}\sim\mathrm{N}(20,50^{2})$; $X_{4i}\sim\mathrm{N}(50,50^{2})$; $X_{5i}\sim\mathrm{N}(2,5^{2})$ and $X_{6i}\sim\mathrm{N}(25,50^{2})$.
@@
    Define the functional coefficients $\beta_{0}(t)=\exp(t^{2})$; $\beta_{3}(t)=\cos(2t)$ and $\beta_{5}(t)=t^{3}$, in which $t\in[0,2]$;
@@
    All other partial functional coefficients do not need to be defined since it is considered that only two covariates $X_{3i}$ and $X_{5i}$ should be used for data generation;
@@
The equality presented in (5) is a summary representation, since its full version is given by
@@
The covariates chosen to compose the discrete representation (5) were purposely positioned in the design matrix $\boldsymbol{X}$ so that both were not neighbors. Thus, we can verify that the model truly selects covariates, not just blocks of covariates.
Therefore, any method that provides a good fit to this dataset should be able to identify that only $X_{3i}$ and $X_{5i}$ are useful covariates, and yield estimates of the functional coefficients that are close to the original functional parameters. When building the synthetic data, we establish that $\beta_{3}(t)$ and $\beta_{5}(t)$ are the so-called “functional coefficients” instead of “partial functional coefficients” since it is implied that the latent variable is equal to one in the generation of synthetic data in both cases.
@@
Bayesian inference was implemented through the Gibbs sampler described in Subsection 2.1 using two chains to enable the convergence diagnosis, which start at different points, and with the calculation of 10 000 iterations. Considering a burn-in period of 50% of the size of each chain and spacing of 50 points between each sampled value, each chain contains 100 sampled points. Thus, at the end of the process, a posterior sample of size 200 is obtained for each parameter and latent variable.
The convergence diagnosis used in our analyses takes into account the convergence test of the chains proposed by Gelman and Rubin (1992). In our numerical experiments carried out herein, the chains were initialized as follows:
@@
The latent random variables $Z_{l}$’s were initialized in the first chain via random draw, and the complementary vector of the drawn variables for the first chain was used for the second chain. In situations where $\boldsymbol{\mu}$ is considered as a parameter, the first chain was initialized with $\boldsymbol{\mu}=\boldsymbol{\frac{1}{5}}$, while the second was with $\boldsymbol{\mu}=\boldsymbol{\frac{4}{5}}$.
To summarize the results, we calculated, for each parameter, the MAP (Maximum a Posteriori) estimate. Thus, regarding the latent variables $Z_{l}$’s, we obtained the most frequent value, meaning the posterior mode.
@@ 3.3 Results
Supplementary Figures 1 through 4 summarize the boxplots of the metrics (7) and MSE across all tested model configurations in our simulations, including scenarios in which $\mu$ is treated as a hyperparameter or a parameter. The results consistently demonstrate that the proposed model exhibits low sensitivity to the choice of $\mu$, regardless of the dispersion level of the data or how $\mu$ is treated. This robustness ensures the proposed model can effectively smooth and select partial functional coefficients (and, consequently, covariates) regardless of the prior choice for $\mu$. Both metrics exhibit expected behaviour concerning $\sigma$, with improved performance when data are generated with low dispersion ($\sigma=0.2$). Additionally, the symmetry and limited number of outliers in the boxplots confirm the well-behaved nature of the metrics, highlighting the model’s efficiency and reliability across configurations.
As an example of results from one of the simulated datasets, Figure 1 shows the estimates $\hat{\beta}_{l}(t)$ of the partial functional coefficients selected by the model, along with 95% credible bands for a configuration with $\mu=0.1$ under two dispersion levels ($\sigma=0.2$ and $\sigma=20$). The credible bands were computed using two hundred parameter vectors sampled by the Gibbs sampler, generating two hundred corresponding curves for each partial functional coefficient. Quantiles at $2.5\%$ and $97.5\%$ were then used to define the lower and upper bounds at each point $t$.
@@
We can also observe in Figure 1 that the proposed model demonstrates strong fitting performance, particularly with low dispersion data, where the true and estimated curves overlap almost perfectly. Although precision diminishes with higher noise levels, the estimates remain accurate. Notably, the model assigns near-zero estimates to the partial functional coefficients of excluded covariates, as shown in Supplementary Figure 5. This pattern reflects the dependence of the distributions of the basis expansion coefficients on the respective latent variables. Additionally, Figure 2 presents the first three estimated mean (or expected) functional response curves (among the total of $m=10$) alongside their credible bands, true curves, and observed data. The credible bands, constructed similarly to those in Figure 1, further demonstrate the model’s ability to capture mean responses across different dispersion levels effectively.
@@
In addition to goodness-of-fit, it is also interesting to assess the variable selection performance of each model configuration across the one hundred replications. Table 1 presents the proportion among the 100 replications that each functional coefficient estimate ($\hat{Z}_{l}\hat{\beta}_{l}(.)$, for $l=1,\ldots,6$) was different from zero. For data with a low dispersion level ($\sigma=0.2$), we can observe in Table 1 that all tested model configurations correctly performed the selection in all replications. However, the selection of data with higher dispersion ($\sigma=20$) is not perfect in some replications, although it remains highly accurate.
@@
Table 1 shows that except for configurations with $\mu$ as a parameter and as hyperparameter with $\mu=0.1$ when $\sigma=20$, all other configurations (for both $\sigma=0.2$ and $\sigma=20$) correctly selected the covariates associated with the partial functional coefficients $\hat{\beta}_{3}(.)$ and $\hat{\beta}_{5}(.)$ in all replications, allowing us to conclude that the model has no difficulty in selecting the partial functional coefficients that should remain in the functional regression. Furthermore, we can see that most of the selection errors committed in the scenario with $\sigma=20$ happened when the model failed to exclude a coefficient that should have been excluded. The proportions associated with the coefficients that should be excluded from the regression are generally below 7%. However, it is worth noting that slightly higher proportions were observed for the configuration that considers $\mu=0.9$.
@@
Similarly, Supplementary Figure 6 shows the estimated partial functional coefficients ($\hat{\beta}_{3}(.)$ and $\hat{\beta}_{5}(.)$) and functional intercept for the case when the model considers $\mu$ as a parameter. The conclusions are similar to those obtained from Figure 3, showing that the version with more layers in the hierarchical model does not differ significantly from the version that considers $\mu$ fixed.
To complement the analysis of the proposed methodology through simulations, Supplementary Figure 7 presents the estimates of all the other partial functional coefficients ($\hat{\beta}_{1}(.)$, $\hat{\beta}_{2}(.)$, $\hat{\beta}_{4}(.)$ and $\hat{\beta}_{6}(.)$) from the model that considers $\mu$ as a hyperparameter that are associated with the excluded variables, according to the data dispersion degree (Supplementary Figure 5 presents the corresponding plots obtained by the model that considers $\mu$ as a parameter). Regardless of the data dispersion degree and the approach to $\mu$, Supplementary Figures 5 and 7 show that all the functional estimates in question are around zero.
3.3.1 Comparative analysis with grLASSO, grSCAD, grMCP and BGLSS